Blog

Sound studio: Harpex plugin update

A newer version of the Harpex plugin for decoding of ambisonics B-format signals was installed at the sound studio today. I've installed the 64-bit bit version of the plugin (VST and AudioUnit). If anyone need the 32-bit version, please let me know.

0 comments

Possible updates of router over next few days

Uninett has informed us that they might be updating routers over the next few days. This might cause one or more short periods where network to and from BEK (and BEK servers) might be unavailable.

I forbindelse med en nylig publisert sårbarhet i JUNOS, ser vi oss nødt
til å foreta en hasteoppgradering av endel av våre rutere.

Oppgraderingern kan føre til noen små kortvarige forstyrrelser i perioden.

Arbeidet vil starte fra idag klokken 15:00 og vil pågå noen dager
fremover. Vi vil komme med informasjon når arbeidet er utført.

For kjerne-ruterne vil vi gjøre vårt beste for at forstyrrelsene av
trafikken skal bli minimal, siden vi der har alternative veier tilgjengelig.

For mindre rutere ute hos institusjonene vil institusjonene som er
tilkoblet vedkommende ruter oppleve to-tre korte brudd, hver på under
5 minutter mens oppgraderinga blir gjort.

Vi beklager den korte varslinga, men dette er dessverre en nødvendig
oppgradering det haster med.

0 comments

Mailserver down brielfly this morning

The mailserver was down for a short period this morning.

0 comments

Maintenance on mail server

The mail server have been done for maintenance today. We have now fixed the spam filters, and this should reduce the amount of spam received quite a bit as compared to the last few weeks.

0 comments

Firewall updated

The hard drive on the firewall gave up life yesterday. The firewall was up and running again approx 01:30 tonight. We expect delayed mails to arrive over the next few hours.

0 comments

Old server (Odin) has been shut down

Our old web and mail server, Plot, has been shut down. All services were moved to a new server (Odin) last summer. we do not expect this to cause problems for any of our users, but if you do experience difficulties, please contact us.

0 comments

The mail server now use SPF

We are now using SPF (Sender Policy Framework) on the mail server to reduce the amount of spam as well as prevent spambots from pretending that spam mails are sent from bek.no mail accounts.

0 comments

Mail server down

The mail server is unfortunately down at the moment, due to problems caused by a OSX server security update. We are working on it, but it will probably take time do get sorted out again.

Incoming mails are queued on the server, and should be delivered once we get this sorted out again.

With the exception of mail and mailing lists, all other services seems to be working properly.

We apologies for the inconvenience.

0 comments

New web pages

We are in the process of rolling out the new web pages for BEK. While doing so, we observe some glitches that we are looking into.

The following domains need to have DNS updated. This must be done by the domain owners:
  • closterbomb.no
  • eyeofthestorm.no
  • kate.no
  • logografer.no
  • online-printhouse.com
  • www.prettysister.com
  • www.stormensauge.no
  • www.sustainablesource.net/


There are problems with the following domains that we are looking into:

0 comments

Server restarted today

The server was restarted a few minutes ago due to work on collaborative web services.

0 comments

Work on web services

Due to ongoing work on web services, the web services might be down for a few minutes from time to time today.

0 comments